OpenStudy (anonymous):

e^(2sinx)=1 ; 0<=x<=16

OpenStudy (anonymous):

the only way for \[e^y=0\] is if \[y=0\] so you should solve \[2\sin(x)=0\] or \[\sin(x)=0\]

myininaya (myininaya):

this one was pretty :( you stole it satellite

OpenStudy (anonymous):

lots of choices on the interval [0,16] but i will let myininaya write them for you

OpenStudy (anonymous):

take the ln of both sides \[\ln (e ^{2Sinx})=\ln(1)\]

myininaya (myininaya):

5pi is approximately 16 so we are basically looking at the interval [0,5pi]

OpenStudy (anonymous):

then for logarithmiic rules you get \[2\sin (x)=0\]

myininaya (myininaya):

sinx is 0 when x=0,pi,2pi,3pi,4pi,5pi

myininaya (myininaya):

remember 5pi<16 so we include 5pi

OpenStudy (anonymous):

2sin(x)=0 this equation is cero when x=0 or x=180 son the only answers would be 0 if the parameter is in radians the woudl be 0, pi, 2pi , 3pie 4pie ..............5,16pie

myininaya (myininaya):

can i have some pie?